MOLECULAR BEAM STUDY of AlND_{3} COMPLEX: STRUCTURE, VIBRATIONAL FREQUENCIES, and IONIZATION POTENTIAL.

\underline{ZYGMUNT J. JAKUBEK} and BENOIT SIMARD, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON, Canada K1A 0R6.

AlND_{3} complexes are produced in a Smalley-type laser ablation source and probed by resonance 2-photon ionization technique. In the range 18500-23500 cm^{−1}, several vibronic bands for five isotopomers, Al^{14}NH_{3}, Al^{15}NH_{3}, Al^{14}ND_{3}, Al^{14}NHD_{2}, and Al^{14}NH_{2}D, are observed. Only the AlND_{3} complex is studied in detail. The spectrum is assigned to the B̃^{2}A_{1}–X̃^{2}E electronic transition, with the B̃^{2}A_{1} and X̃^{2}E states correlating to the Al(4s)+ND_{3}(^{1}A′_{1}) and Al(3p)+ND_{3}(^{1}A′_{1}) asymptotes, respectively. The AlND_{3} complex has approximate C_{3v} symmetry in both states. The origin band of the B̃^{2}A_{1}–X̃^{2}E transition is located at 18532.5±1.0 cm^{−1}. The spin-orbit splitting of the ground state is determined to be 55.7±0.4 cm^{−1}. Frequencies of the ν_{3} (Al-ND_{3} stretch), ν_{4} (AlN-D_{3} symmetric stretch), and ν_{6} (Al-ND_{3} bend) vibrations in the B̃^{2}A_{1} state are measured to be 315.7 ± 0.6 cm^{−1}, 2657±1 cm^{−1}, and 411±1 cm^{−1}, respectively. Partially resolved rotational structure of some of the bands is observed and analyzed. The adiabatic ionization potential of the AlND_{3} complex is measured by photoionization efficiency technique as equal 39710±10 cm^{−1}. Rydberg series converging to various levels of the Al^{+}-ND_{3} stretching mode of the AlND_{3}^{+} cation are observed. The ν_{3}^{+} frequency is determined to be 322 cm^{−1}. The structure determination of the neutral and monopositive complex and assignment of vibrational modes are aided by ab initio calculation.